I have 345 mL of a 2.5 M NaCl solution. If I boil the ater until the volume of the solution is 250 mL, what will the molarity of the solution be?
1
Expert's answer
2020-04-24T13:35:47-0400

For the beginning we should calculate the amount of substance of sodium chloride:

"\nu(NaCl) = C•V" = 2.5(M)•0.345(L) = 0.85 mol

Since the amount of substance during evaporation doesn't change, so the new concentration will be calculated:

"C(2) = \nu/V(2)" = 0.85(mol)/0.245(L) = 3.4 M

Answer: C(2) = 3.4 M